# Improving Similar Case Retrieval Ranking Performance By Revisiting RankSVM

### v1

@@ 3.1 Results on LeCaRDv1
Model NDCG@10 NDCG@20 NDCG@30
BERT 0.7896 0.8389 0.9113
BERT+RankSVM 0.7963 0.8504 0.9166
BERT+LEVEN 0.7881 0.8435 0.9136
BERT+LEVEN+RankSVM 0.7931 0.8452 0.9164
MAMBA 0.7469 0.8013 0.892
Lawformer 0.7592 0.8169 0.8993
Lawformer+RankSVM 0.7738 0.8258 0.9073
@@
(b) Visualization of BERT Features
(c) Visualization of Rank-SVM Scores
@@
What’s more, the mechanism of RankSVM is actually maximizing ROC(Receiver Operating Characteristic curve)([1]). So we use the model of BERT comparing the results with and without RankSVM to show how ROC changes. For the BERT model, we use the first fold and the C is 0.01, and the visualization result is shown in Figure 7. It could be seen that RankSVM helps improve the result of BERT on the AUC score, and it does an even better job in the test dataset. It proves our guess before that it could help reduce overfitting.
(a) ROC Curve for LeCaRDv1 Training Dataset
@@
Dataset
Figure 7: Three simple graphs
Additionally, to explore SVM’s performance on longer text for retrieval, we also benchmark its performance using Lawformer. It could also be seen that all NDCG-related metrics get improved.
We also conduct experiments using MAMBA. As the performance is not competent enough, we don’t do further combinations with RankSVM.
3.2 Results on LeCaRDv2
The result is shown in Table 4.According to the table, we could find that the NDCG related metrics are generally improved similar to LeCaRDv1, except for the BERT model. To understand its performance, here we also do visualization using features extracted from the BERT model following the same steps as the LeCaRDv1 dataset, see Figure 8 and 9.
@@ 3.2 Results on LeCaRDv2
It is shown by the figure with real labels that BERT features fail to distinguish these two classes explicitly this time as there are a lot of overlaps between the two classes compared with what is extracted from the LeCaRDv1 dataset, while RankSVM tries to learn from those features that are not distinguished between these two classes continuously. The failure of BERT may be owing to the changed standards of relevance in LeCaRDv2 and the shorter document length of the LeCaRDv2 dataset. It explains why RankSVM added to BERT fails to help improve NDCG metrics here.
4 Discussion
4.1 Impact of Length of Text
Even after we changed the length of the input data from 512 to 800 using the MAMBA model and Lawformer model, we could see that the interesting thing here is that there is no apparent increase in metrics compared with BERT, which is also mentioned in other works([6]).
But we could also argue that there lies some space for hyperparameter tuning in the future.
4.2 Error Analysis
To explore further why RankSVM fails to improve the performance on BERT-based models regarding the LeCaRDv2 dataset, we conduct experiments using multi-class BERT classification to extract features out of our speculation that RankSVM on LeCaRDv2 needs language models that provide more information and end up getting all NDCG-related metrics improved again, see table 5. So our speculation holds as RankSVM still works for BERT on the LeCaRDv2 dataset under this setting. Recently, there have been more research studies on improving RankSVM, and we leave that to explore in the future.
@@ 4.2 Error Analysis
5 Conclusion
In our paper, we examine the performance of RankSVM when combined with other language models using binary labels to serve a similar case retrieval task on the LeCardv1 dataset and the LeCardv2 dataset. We come to the conclusion that RankSVM with binary labels could help improve the performance of models by improving their concrete ranks, and we also try to give some explanation from the feature perspective. Also, the RankSVM could help fight overfitting, which results from an imbalance class common in this task. Our findings point to the potential of improving SCR task performance from an optimization perspective. There still lies some work that needs to be done to explore the appropriate model to improve performance on the recall metric on the LeCaRDv2 dataset and how the performance of RankSVM will change when we extract features with larger batch sizes. We also leave the investigation regarding RankSVM combined with other models and datasets for future work.

### v2

@@ 4.2 Experiments on LeCaRDv1
Model NDCG@10 NDCG@20 NDCG@30 P@5
BERT 0.7896 0.8389 0.9113 0.4556
BERT+RankSVM 0.7963 0.8504 0.9166 0.4571
BERT+LEVEN 0.7881 0.8435 0.9136 0.4637
BERT+LEVEN+RankSVM 0.7931 0.8452 0.9164 0.4717
MAMBA 0.7136 0.7760 0.8744 0.3888
MAMBA+RankSVM 0.7144 0.7743 0.8742 0.4002
Lawformer 0.7592 0.8169 0.8993 0.4081
Lawformer+RankSVM 0.7738 0.8258 0.9073 0.4033
From the results, we could see that RankSVM helps improve performance on NDCG and has exceptional performance on precision for the top 5. And the advantage in NDCG is more obvious for the more relevant cases, which proves that our model is better at putting highly relevant cases earlier in a similar case list. To better understand the advantage of RankSVM, we finish visualization based on the training and testing dataset on fold 0 to see the score computed by original BERT and BERT+RankSVM.
Specifically, we use features extracted by BERT as our input and compress them into two dimensions using t-SNE[30]. After that, we color them according to similarity scores computed by relative models or labels. See Figure 5 and 6 for reference.
@@
(b) Visualization of BERT Scores
(c) Visualization of RankSVM Scores
@@
What’s more, the mechanism of RankSVM is actually maximizing ROC(Receiver Operating Characteristic curve) pointed out in some early paper [31], said to outperform linear SVMs, perceptron and logit. The AUC is especially suitable for class imbalance setting. So we use BERT with RankSVM and BERT without RankSVM to compare their ROC and AUC values. We use the first fold data and the C is 0.01. The visualization result is shown in Figure 7 and 8. It could be seen that RankSVM helps improve the result of the BERT model on the AUC score, and it does even better in the test dataset. It proves our guess before that it could help reduce overfitting without sacrificing efficiency.
(a) ROC Curve for LeCaRDv1 Training Dataset with BERT based models
@@
Dataset with BERT based models
Figure 7: ROC Curves Using BERT model and BERT+RankSVM
(a) ROC Curve for LeCaRDv1 Training Dataset with MAMBA based models
(b) ROC Curve for LeCaRDv1 Test
Dataset with MAMBA based models
Figure 8: ROC Curves Using MAMBA model and MAMBA+RankSVM
4.3 Experiments on LeCaRDv2
The result on LeCaRDv2 is shown in Table 4. As the performance of MAMBA is not as competitive in LeCaRDv1, we don’t use it for LeCaRDv2. According to the table, we could find that the NDCG related metrics are generally improved similar to LeCaRDv1, except for the BERT model.
To understand its performance, here we also do visualization using features extracted from the BERT model following the same steps as the LeCaRDv1 dataset, see Figure 9 and 10.
@@ 4.3 Experiments on LeCaRDv2
It is shown by the figure with real labels that BERT features fail to distinguish these two classes explicitly this time as there are a lot of overlaps between the two classes compared with what is extracted from the LeCaRDv1 dataset, while linear RankSVM tries to learn from those features that are not distinguished between these two classes continuously. The failure of BERT may be owing to the changed standards of relevance in LeCaRDv2 and the shorter document length of the LeCaRDv2 dataset. The visualization explains why RankSVM added to BERT fails to help improve NDCG metrics here.
4.4 Comparison Study
To further explore the features of RankSVM and the potential of pairwise ranking methods.We finish comparison experiments with RankNet using some fraction of data from LeCaRDv1, and the result is shown as follows, see table 5. We can tell that RankNet is also very competitive especially considering its precision and it proves that there is a lot to explore with pairwise methods.
Methods NDCG@10 NDCG@20 NDCG@30 P5 P10
RankSVM 0.6995 0.7711 0.8576 0.3 0.29
RankNet 0.7816 0.8336 0.8432 0.5 0.62
Table 5: Results on Comparison Experiments with RankNet
5 Conclusion
Even after we changed the length of the input data from 512 to 800 using the MAMBA model and Lawformer model, we could see that there is no apparent increase in metrics compared with BERT, which is also mentioned in other works[32]. It is worth exploring how long-input language models could really benefit downstream applications.
To investigate the reason why RankSVM fails to improve the performance on BERT-based models on the LeCaRDv2 dataset, we conduct experiments using multi-class BERT classification to extract features. We speculate that RankSVM on LeCaRDv2 needs language models that provide more information for ranking as Lawformer and BERT+LEVEN model have provided more information. We end up getting all NDCG-related metrics improved again, see table 6. RankSVM still works for BERT on the LeCaRDv2 dataset under this setting.
@@ 5 Conclusion
In our paper, we examine the performance of RankSVM when combined with other language models using binary labels on SCR tasks on the LeCardv1 dataset and the LeCardv2 dataset. We come to the conclusion that it could help improve the performance of models by improving their concrete ranks, and we also give explanations from the feature extraction perspective. The improvements are more obvious for more relevant documents. Also, the RankSVM could help mitigate overfitting with surprising efficiency, which results from an imbalance class common in this task. Our findings point to the potential of improving SCR task performance from a learning-to-rank perspective, especially considering pairwise methods.
While previous work has considered different ranking functions, there has been no research focus on their difference and their suitability for SCR tasks, especially pairwise methods. There still lies some work that needs to be done to explore the appropriate ranking function to improve performance on the recall metric on the LeCaRDv2 dataset and how to exploit long inputs to improve performance. We also leave the investigation regarding other pairwise methods combined with other models and datasets for future work.
Declaration on Generative AI
During the preparation of this work, Yuqi Liu used the GPT model in order to: Grammar and spelling check. After using these tool(s)/service(s), the author(s) reviewed and edited the content as needed and take(s) full responsibility for the publication’s content.